Let Z be a standard normal random variable. Use the calculator provided, or this table, to determine the value of c.

P(Z <_c) = 0.2061

Carry your intermediate computations to at least four decimal places. Round your answer to two decimal places.
Answer #1

Solution :

=> P(Z <= c) = 0.2061

=> c = -0.8200

=> P(Z <= -0.8200) = 1 - P(Z <= 0.8200)

= 1 - 0.7939

= 0.2061